# Technology, Institution, and Regional Growth: Evidence from Mineral Mining Industry in Industrializing Japan

### v14

@@ 4.3 Dependent Variables
The primary dependent variable, population, was obtained by digitizing the municipal-level statistics documented in the 1930 and 1935 Population Censuses (Statistics Bureau of the Cabinet 1935, 1939).
@@ 5.1 Estimation Strategy
$\footnotesize{\begin{split}y_{i}=\alpha+\beta\text{{MineDeposit}}_{i}+\mathbf{x}_{i}^{\prime}\bm{\gamma}+e_{i},\end{split}}$ (2)
where $i$ indexes municipalities, $y$ is the outcome variable, MineDeposit is an indicator variable that equals one for municipalities within 5 km from a mine, $\mathbf{x}$ is a vector of control variables, and $e$ is a random error term. As the placement of a mine is determined by a geological anomaly, MineDeposit is random in nature (Benshaul-Tolonen 2019, p.1568). However, the rest of the variation may be correlated with the local variation in infrastructure. Thus, if a mineral deposit is found between a village and a city, the mining firm may choose the city as its main mining point because the city is likely to have better infrastructure than the village. Then, MineDeposit can be positively correlated with the error term, leading to a positive (negative) omitted variable bias in the estimate of $\beta$ if the placement of the city is positively (negatively) correlated with the outcome variable.333333 As an example, consider a simplified projection of $y$ on MineDeposit and City, $y=\alpha+\beta\text{{MineDeposit}}+\gamma\textit{City}+e$. When the municipality type (City) is omitted, the linear projection coefficient can be written as $\beta^{*}=\beta+\Xi\gamma$, where $\Xi=(E[\text{{MineDeposit}}^{2}]^{-1})(E[\text{{MineDeposit}}\text{{City}})$. As explained, MineDeposit and City may be positively correlated such that $\Xi>0$. In addition, when $y$ is the population, conditional on mine deposits, the city has a greater number of people than town/villages ($\gamma>0$). This implies that $\beta^{*}=\beta+\Xi\gamma>\beta$. To deal with this systematic bias in the ordinary least squares, two indicator variables are included that equal one for cities and towns, respectively. As explained in Section 4.4, the distances to the nearest neighboring railway station, seaport, and river and the elevation are included to explicitly control for transportation accessibility. [^34] The differences in the estimates from both specifications, including and excluding the control variables, indicate how much this omitted variable mechanism influences the results. Thus, the randomness of the primary exposure variable can be partially assessed (MineDeposit). As demonstrated later, the results from the specification including the city and town fixed effects are materially similar to those from the specifications. This includes the control variables, thereby supporting the randomness of the exposure variable. To be conservative, the specification including the control variables is preferred.
@@
$\footnotesize{\begin{split}\textit{MineDeposit}_{i}=\kappa+\tau\textit{Stratum}_{i}+\mathbf{x}^{\prime}_{i}\bm{\mu}+\epsilon_{i},\end{split}}$ (3)
@@
The heteroskedasticity-consistent covariance matrix estimator is used as a baseline estimator. [^36] The results are materially similar between the Eicker-White type covariance matrix estimator suggested by Hinkley (1977) and the HC2 estimator proposed by Horn et al. (1975). For the reduced form regressions, to be conservative, the HC2 estimator is used given that the treatment variable is relatively sparse (Section 4.1). For the sensitivity check, the standard errors clustered at the county level based on the cluster-robust covariance matrix estimator (Arellano 1987) are used to determine the influence of the potential influences of the local-scale spatial correlations. Online Appendix D.3 shows that the results are materially similar under both variance estimators. Therefore, the potential spatial correlations were negligible in this empirical setting.
@@ 5.2 Main Results
Panel A of Table 2 presents the main results for the population in 1930. The estimate from a simple IV estimation is shown in column 1.373737The first-stage $F$-statistics reported in the same panel exceed the rule-of-thumb threshold value proposed by Staiger and Stock (1997), say 10, in all regressions. This shows that the necessary rank condition is satisfied. Online Appendix D.3 confirms that the first-stage $F$-statistics never break this threshold when a conservative variance estimator such as the cluster-robust variance-covariance matrix estimator is used. Then I include the city and town fixed effects (column 2), and both the fixed effects and control variables (column 3). The estimated coefficient decreases from $1.386$ to $0.829$ when the fixed effects are included. This suggests that, as expected, the location of the coal mines may be positively correlated with the local infrastructure. Including additional control variables on transportation accessibility has a moderate influence on the estimate ($0.682$). This supports the evidence that, after conditioning on the fixed effects, there are few influential unobservables that are potentially correlated with the location of mines. Panel B confirms the similar results for the population in 1935 in the same column layouts. The estimated coefficient from the specification including the full set of controls in column 3 is $0.723$, which is larger than that for the 1930 sample.
The estimated magnitude is economically meaningful. The estimates from the preferred specification of column (3) indicate that coal mines increased the local population by $98$% [^38] Statistically, it is more precise to state that the geometric mean of population in the treatment group with mines was approximately 98% (i.e., $\text{exp}(0.682)=\text{exp}(\hat{\alpha}+0.682)/\text{exp}(\hat{\alpha})\approx 1.98$) greater than that of the control group. However, simple interpretations are used throughout this study to avoid redundancy. in 1930 (panel A) and $106$% in 1935 (panel B). The average number of miners per coal mine increased just less than $100$ during this period ($1,567$ in 1931 and $1,658$ in 1936). [^39] This is because while the number of male miners increased, simultaneously the number of female miners decreased due to the revision of the regulation (Table 1). From this perspective, it is surprising that the magnitude is estimated to be greater in the 1935 sample. The mechanism behind this trend is assessed in the next subsections.
@@ 5.3 Fertility Responses
The results show that marriage and fertility rates have reverted to their natural trends after the regulation was enforced. This aligns with the prediction described in Section 4, supporting the evidence that the declines in female relative wages had shifted women to take on the role of family building. [^41] The optimal ratio of home labor is given by $\frac{h_{m}}{h_{f}}=\left(\frac{\kappa_{m}}{\kappa_{f}}\frac{w_{f}}{w_{m}}\right)^{\frac{1}{\rho}}$ (Online Appendix B). The labor force participation rate is only available in the 1930 Population Census. Online Appendix D.4 confirms that the mining area had lower female labor force participation than the surrounding agrarian area, with a structural shift from the agricultural to the mining sector. This is consistent with the fact that the mining area has lower female relative wages than the agrarian area (Section 2.2), supporting the labor supply mechanism behind the fertility choice. Although the data on the labor force participation rate are unavailable in the 1935 Population Census, the time-series statistics also confirm that the revision of the regulations decreased the number of female miners (Section 2.2). Under limited substitutions with the husband’s home labor and purchased services, the increase in the gender wage gap pushed the females out of the local labor market and shifted them to domestic work. This forced women to take on the role of building a family.
@@ 5.4 Mortality Changes
Turning to look at columns 3–4, the estimated coefficient for the female mortality rate alone is negative and weakly statistically significant. The estimate is $-2.010$, roughly half of the standard deviation of female mortality in 1935. Thus, this may support the former occupational hazard channel. The institutional change from the enforcement of regulation liberated women from dangerous underground work, which may have improved the overall health status of females.
@@
The result for overall mortality suggests that the regulation mitigated the female miners’ occupational hazards. To illustrate a more comprehensive view of the improvements in local mortality risk, the extent to which occupational hazards influenced the reproductive health of female miners and whether institutional change improved such hazards are assessed. This is done using the early-life mortality measured in the registration-based vital statistics. The potential influence of mining pollution in the mining area is also assessed.
@@
First of all, a theoretical framework is illustrated for mortality selection in utero to interpret the quantitative evidence from the vital statistics data. The biological sorting mechanism in utero suggested by the Trivers-Willard hypothesis (TWH) (Trivers and Willard 1973) implies that there are two possible cases in which fetuses are exposed to health shocks: scarring and selection.434343 Valente (2015) tests the TWH using the case of the civil conflict in Nepal and shows that fetal exposure to the war decreased the sex ratio at births.
@@
Notes: Panels A and B show the results for the 1933 and 1938 Vital Statistics samples, respectively. The IMR, FDR, and CDR indicate the infant, fetal, and child mortality rates, respectively. Columns 1–2 show the results for the entire sample. Columns 3–4 show the results for the municipalities with coal mines (i.e., the municipalities included in the treatment group). Columns 5–6 show the results for the full sample and the municipalities with coal mines, respectively. Panel B of Table C.4 summarizes the summary statistics of the early-life mortality variables by treatment and control groups.
@@
Next, we test whether the occupational hazard led to this scarring due to coal mining. In column 3, the sample is limited to $93$ municipalities with coal mines to analyze the correlation between the FDR and the number of female miners and accessibility to rivers. If heavy manual work during pregnancy affects fetal health, FDRs should be positively correlated with the number of female miners. To control for the scale effect and provide a placebo test, the number of male workers was included in the same specification. The estimated coefficient for female miners is significantly positive, whereas that for male miners is negative. This result supports the evidence that occupational hazards for females (not males) alone increased the risk of death before birth. The estimate indicates that a one standard deviation increase in female miners ($=269.2$ miners) increases the number of fetal deaths by roughly $10$ ($0.038\times 269$) per 1000 births. This is economically meaningful because the mean difference in the FDR between the treatment and control groups was also approximately $9$ fetal deaths per 1000 births (Panel B of Online Appendix Table C.4). Column 4 shows the result for the IMR, showing that the estimated coefficient on the number of female miners is also statistically significantly positive. It further provides evidence that this reduction in fetal health endowment via occupational hazards for females is associated with higher infant mortality risks.
@@
Panel B of Table 5 provides the results for the 1938 sample in the same panel and column layout. Columns 1–2 show positive but much smaller estimates, which are no longer statistically significant. The estimates listed in columns 3–4 are also positive, but smaller in magnitude. The estimate for the FDR shows that a one standard deviation increase in female miners ($=179.3$ miners) increased the number of fetal deaths by approximately $6$ ($0.033\times 179$) per 1000 births. This magnitude was roughly $60$% of that in 1933.
@@
The decline in the estimated magnitude of these health outcomes between 1933 and 1938 is noteworthy. Although this could partly be associated with an improving trend in health-related risk-coping strategies such as increments in the number of medical doctors and installation of modern water supply (Online Appendix Figure D.5), the city and town fixed effects and accessibility variables can control for the heterogeneities in the local sanitary levels. Therefore, my results imply that the occupational hazards had increased the early-life mortality risk through the scarring mechanism in utero. However, the risks substantially decreased in the late 1930s because enforcing the revised labor regulations changed the labor patterns of females and reduced the number of female miners. Consequently, the institutional change may have attenuated the mortality selection mechanism before births throughout the 1930s.
@@
Section 2 suggests that the coal sludge generated from the coal selection process could have been a potential pollution factor. If such wastewater increases the health risk of fetuses and infants, municipalities with mines closer to rivers should have significantly higher mortality rates because the wastewater had been discharged into the rivers (Chiba and Yamada 1964). However, the estimated coefficients of the river accessibility variable are close to zero and statistically insignificant in columns 3 and 4 in both panels. This result suggests that, while wastewater was occurring, it was not suffincient to harm the health conditions of mothers and infants, as anecdotes suggest (Section 2.3).
Moreover, air pollution is also unlikely to increase mortality risks. If air pollution increases early-life mortality, then the estimated coefficients on the male miner should be positive in columns 3 and 4. This is because larger mines should emit more coal smoke from the boilers of the steam winding machines. As shown, however, the estimates are very close to zero in all the cases. Column 5 of Panel A further tests the potential influence of air pollution using an alternative outcome, the mortality rate of children aged 1–4. If air pollution was a contributory factor, the child mortality rate would be greater in the coal mining area than in the surrounding regions. This is because younger children are more susceptible to pollutants and are more likely to play outside than infants. The estimate is, however, close to zero and is statistically insignificant. Column 6 of Panel A uses data on $93$ municipalities with coal mines and provides further evidence. The estimated coefficients on the number of female and male miners are close to zero, suggesting that child mortality did not depend on the scale of emissions. The results for child mortality in 1938 are materially similar to those in 1933 (columns 5–6 in Panel B). Overall, air pollution does not appear to be a plausible channel to explain the greater early-life mortality in the mining area. This is consistent with the historical fact that winding machines were electrified during the interwar period (Section 2.3).

### v15

@@ 4.3 Dependent Variables
The primary dependent variable, population, was obtained by digitizing the municipal-level statistics documented in the 1930 and 1935 Population Censuses (Statistics Bureau of the Cabinet 1935a, 1939).
@@ 5.1 Estimation Strategy
$\footnotesize{\begin{split}y_{i}=\alpha+\beta\text{{MineDeposit}}_{i}+\mathbf{x}_{i}^{\prime}\bm{\varphi}+e_{i},\end{split}}$ (2)
where $i$ indexes municipalities, $y$ is the outcome variable, MineDeposit is an indicator variable that equals one for municipalities within 5 km from a mine, $\mathbf{x}$ is a vector of control variables, and $e$ is a random error term. As the placement of a mine is determined by a geological anomaly, MineDeposit is random in nature (Benshaul-Tolonen 2019, p.1568). However, the rest of the variation may be correlated with the local variation in infrastructure. Thus, if a mineral deposit is found between a village and a city, the mining firm may choose the city as its main mining point because the city is likely to have better infrastructure than the village. Then, MineDeposit can be positively correlated with the error term, leading to a positive (negative) omitted variable bias in the estimate of $\beta$ if the placement of the city is positively (negatively) correlated with the outcome variable.353535 As an example, consider a simplified projection of $y$ on MineDeposit and City, $y=\alpha+\beta\text{{MineDeposit}}+\varphi\textit{City}+e$. When the municipality type (City) is omitted, the linear projection coefficient can be written as $\beta^{*}=\beta+\Xi\varphi$, where $\Xi=(E[\text{{MineDeposit}}^{2}]^{-1})(E[\text{{MineDeposit}}\text{{City}})$. As explained, MineDeposit and City may be positively correlated such that $\Xi>0$. In addition, when $y$ is the population, conditional on mine deposits, the city has a greater number of people than town/villages ($\varphi>0$). This implies that $\beta^{*}=\beta+\Xi\varphi>\beta$. To deal with this systematic bias in the ordinary least squares, two indicator variables are included that equal one for cities and towns, respectively. As explained in Section 4.4, the distances to the nearest neighboring railway station, seaport, and river and the elevation are included to explicitly control for transportation accessibility. [^36] The differences in the estimates from both specifications, including and excluding the control variables, indicate how much this omitted variable mechanism influences the results. Thus, the randomness of the primary exposure variable can be partially assessed (MineDeposit). As demonstrated later, the results from the specification including the city and town fixed effects are materially similar to those from the specifications. This includes the control variables, thereby supporting the randomness of the exposure variable. To be conservative, the specification including the control variables is preferred.
@@
$\footnotesize{\begin{split}\textit{MineDeposit}_{i}=\iota+\zeta\textit{Stratum}_{i}+\mathbf{x}^{\prime}_{i}\bm{\xi}+\epsilon_{i},\end{split}}$ (3)
@@
The heteroskedasticity-consistent covariance matrix estimator is used as a baseline estimator. [^38] The results are materially similar between the Eicker-White type covariance matrix estimator suggested by Hinkley (1977) and the HC2 estimator proposed by Horn et al. (1975). For the sensitivity check, the standard errors clustered at the county level based on the cluster-robust covariance matrix estimator (Arellano 1987) are used to determine the influence of the potential influences of the local-scale spatial correlations. Online Appendix D.3 shows that the results are materially similar under both variance estimators. Therefore, the potential spatial correlations were negligible in this empirical setting.
I aim to estimate the effects of coal mines on local demographic outcomes in each census year and discuss these effects before and after the regulation was enforced. The regulation included grace periods by 1933 (Section 2). As a result, the estimates using the standard policy evaluation technique are attenuated, because it is not possible to define the pure pretreatment period. Although a rigorous evaluation of the regulation policy is beyond the scope of this study, I have confirmed that the findings from the policy evaluation setting that uses a panel structure align well with my baseline results (Online Appendix D.4).
@@ 5.2 Main Results
Panel A of Table 2 presents the main results for the population in 1930. The estimate from a simple IV estimation is shown in column 1.393939The first-stage $F$-statistics reported in the same panel exceed the rule-of-thumb threshold value proposed by Staiger and Stock (1997), say 10, in all regressions. This shows that the necessary rank condition is satisfied. Online Appendix D.3 confirms that the first-stage $F$-statistics never break this threshold when a conservative variance estimator such as the cluster-robust variance-covariance matrix estimator is used. Then I include the city and town fixed effects (column 2), and both the fixed effects and control variables (column 3). The estimated coefficient decreases from $1.386$ to $0.829$ when the fixed effects are included. This suggests that, as expected, the location of the coal mines may be positively correlated with the local infrastructure. Including additional control variables on transportation accessibility has a moderate influence on the estimate ($0.682$). This supports the evidence that, after conditioning on the fixed effects, there are few influential unobservables that are potentially correlated with the location of mines. Panel B confirms the similar results for the population in 1935 in the same column layouts. The estimated coefficient from the specification including the full set of controls in column 3 is $0.723$, which is larger than that for the 1930 sample.
The estimated magnitude is economically meaningful. The estimates from the preferred specification of column 3 indicate that coal mines increased the local population by $98$% [^40] Statistically, it is more precise to state that the geometric mean of population in the treatment group with mines was approximately 98% (i.e., $\text{exp}(0.682)=\text{exp}(\hat{\alpha}+0.682)/\text{exp}(\hat{\alpha})\approx 1.98$) greater than that of the control group. However, simple interpretations are used throughout this study to avoid redundancy. in 1930 (panel A) and $106$% in 1935 (panel B). The average number of miners per coal mine increased just less than $100$ during this period ($1,567$ in 1931 and $1,658$ in 1936). [^41] This is because while the number of male miners increased, simultaneously the number of female miners decreased due to the revision of the regulation (Table 1). From this perspective, note that the magnitude is estimated to be greater in the 1935 sample. [^42] An alternative empirical specification using a panel structure shows that the difference in magnitude between the 1930 and 1935 census data is statistically significant. See Online Appendix D.4 for details. The mechanism behind this trend is assessed in the next subsections.
@@ 5.3 Fertility Responses
The results show that marriage and fertility rates have reverted to their natural trends after the regulation was enforced. Evidence from an alternative estimation strategy that uses a panel structure supports that the fertility in the coal mining area increased after the enforcement of the regulation (Online Appendix D.4). This aligns with the prediction described in Section 4, supporting the evidence that the declines in female relative wages had shifted women to take on the role of family building. [^44] The optimal ratio of home labor is given by $\frac{h_{m}}{h_{f}}=\left(\frac{\kappa_{m}}{\kappa_{f}}\frac{w_{f}}{w_{m}}\right)^{\frac{1}{\rho}}$ (Online Appendix B). The labor force participation rate is only available in the 1930 Population Census. Online Appendix D.5 confirms that the mining area had lower female labor force participation than the surrounding agrarian area, with a structural shift from the agricultural to the mining sector. This is consistent with the fact that the mining area has lower female relative wages than the agrarian area (Section 2.2), supporting the labor supply mechanism behind the fertility choice. Although the data on the labor force participation rate are unavailable in the 1935 Population Census, the time-series statistics also confirm that the revision of the regulations decreased the number of female miners (Section 2.2). Under limited substitutions with the husband’s home labor and purchased services, the increase in the gender wage gap pushed the females out of the local labor market and shifted them to domestic work. This forced women to take on the role of building a family.
@@ 5.4 Mortality Changes
Turning to look at columns 3–4, the estimated coefficient for the female mortality rate is negative and weakly statistically significant. The estimate is $-2.010$, roughly half of the standard deviation of female mortality in 1935. Although the difference in the estimates for 1930 and 1935 is marginal (columns 2 and 4), this may be in line with the former occupational hazard channel.
@@
Given the result for mortality by sex, I evaluate how occupational hazards affect the reproductive health of female miners by examining early-life mortality data from registration-based vital statistics. In addition, I consider the potential influence of mining pollution in the mining area.
@@
First of all, a theoretical framework is illustrated for mortality selection in utero to interpret the quantitative evidence from the vital statistics data. The biological sorting mechanism in utero suggested by the Trivers-Willard hypothesis (TWH) (Trivers and Willard 1973) implies that there are two possible cases in which fetuses are exposed to health shocks: scarring and selection.464646 Valente (2015) tests the TWH using the case of the civil conflict in Nepal and shows that fetal exposure to the war decreased the sex ratio at births.
@@
Notes: Panels A and B show the results for the 1933 and 1938 Vital Statistics samples, respectively. The IMR, FDR, and CMR indicate the infant, fetal, and child mortality rates, respectively. Columns 1–2 show the results for the entire sample. Columns 3–4 show the results for the municipalities with coal mines (i.e., the municipalities included in the treatment group). Columns 5–6 show the results for the full sample and the municipalities with coal mines, respectively. Panel B of Table C.4 summarizes the summary statistics of the early-life mortality variables by treatment and control groups.
@@
Next, we consider whether the scarring is associated with the occupational hazards of mining activities. In column 3, the sample is limited to $93$ municipalities with coal mines to analyze the correlation between the FDR and the number of female miners and accessibility to rivers. If heavy manual work during pregnancy affects fetal health, FDRs should be positively correlated with the number of female miners. To control for the scale effect and provide a placebo test, the number of male workers was included in the same specification. The estimated coefficient for female miners is significantly positive, whereas that for male miners is negative. This result supports the evidence that occupational hazards for females (not males) alone increased the risk of death before birth. The estimate indicates that a one standard deviation increase in female miners ($=269.2$ miners) increases the number of fetal deaths by roughly $10$ ($0.038\times 269$) per $1,000$ births. This is economically meaningful because the mean difference in the FDR between the treatment and control groups was also approximately $9$ fetal deaths per $1,000$ births (Panel B of Online Appendix Table C.4). Column 4 shows the result for the IMR, showing that the estimated coefficient on the number of female miners is also statistically significantly positive. It further provides evidence that this reduction in fetal health endowment via occupational hazards for females is associated with higher infant mortality risks.
@@
Panel B of Table 5 provides the results for the 1938 sample in the same panel and column layout. Columns 1–2 show positive, but smaller estimates than those for the 1933 sample. [^49] Online appendix D.4 considers the panel data structure in mortality, suggesting that the infant mortality and fetal death rates in coal mining areas may have decreased since the enforcement of the regulation. However, it also indicates that there might be regional heterogeneity in these decreasing trends. The estimates listed in columns 3–4 are also positive, but smaller in magnitude. The estimate for the FDR shows that a one standard deviation increase in female miners ($=179.3$ miners) increased the number of fetal deaths by approximately $6$ ($0.033\times 179$) per $1,000$ births. This magnitude was roughly $60$% of that in 1933.
@@
The decrease in the estimated magnitude of health outcomes between 1933 and 1938 is notable, though it does not represent a significant change. This could partly be associated with an improving trend in health-related risk-coping strategies such as increments in the number of medical doctors and installation of modern water supply (Online Appendix Figure D.5). However, the city and town fixed effects and accessibility variables can control for the heterogeneities in the local sanitary levels. Therefore, my results imply that the occupational hazards had increased the early-life mortality risk through the scarring mechanism in utero.
Overall, the risks decreased in the late 1930s because the enforcement of revised labor regulations changed the labor patterns of females and reduced the number of female miners. My results suggest that this institutional change weakened the mortality selection mechanism before births throughout the 1930s.
@@
Section 2 suggests that the coal sludge generated from the coal selection process could have been a potential pollution factor. If such wastewater increases the health risk of fetuses and infants, municipalities with mines closer to rivers should have significantly higher mortality rates because the wastewater had been discharged into the rivers (Chiba and Yamada 1964). However, the estimated coefficients of the river accessibility variable are close to zero and statistically insignificant in columns 3 and 4 in both panels. This result suggests that, while wastewater was occurring, it was not sufficient to harm the health conditions of mothers and infants, as anecdotes suggest (Section 2.3).
Moreover, air pollution is also unlikely to increase mortality risks. If air pollution increases early-life mortality, then the estimated coefficients on the male miner should be positive in columns 3 and 4. This is because larger mines should emit more coal smoke from the boilers of the steam winding machines. As shown, however, the estimates are very close to zero in all the cases. Column 5 of Panel A further tests the potential influence of air pollution using an alternative outcome, the mortality rate of children aged 1–4. If air pollution was a contributory factor, the child mortality rate would be greater in the coal mining area than in the surrounding regions. This is because younger children are more susceptible to pollutants and are more likely to play outside than infants. The estimate is, however, close to zero and is statistically insignificant. Column 6 of Panel A uses data on $93$ municipalities with coal mines and provides further evidence. The estimated coefficients on the number of female and male miners are close to zero, suggesting that child mortality did not depend on the scale of emissions. The results for child mortality in 1938 are materially similar to those in 1933 (columns 5–6 in Panel B).
In short, air pollution does not appear to be a plausible channel to explain the greater early-life mortality in the mining area. This is consistent with the historical fact that winding machines were electrified during the interwar period (Section 2.3).